OpenStudy (anonymous):

COULD YOU SHOW YOU WORK? 1. x – 3 > –9 2. –2x + 1 ≤ –11 3. 10 < –3x + 1 4. 2(x + 5) > 8x – 8 5. –2(x – 3) ≥ 5 – (x + 3)

OpenStudy (anonymous):

1) x-3>-9 +3 +3 x>-6

OpenStudy (anonymous):

1) add 3 to both sides

OpenStudy (anonymous):

2) subtract 1 from both sides divide by -2 change the sense of the inequality

OpenStudy (anonymous):

satellite73 which on are you doing?

OpenStudy (anonymous):

\[-2x+1\leq -11\] \[-2x\leq -11-1\] \[-2x\leq -12\] \[x\geq -12\div -2\] \[x\geq 6\]

OpenStudy (anonymous):

that as #2

OpenStudy (anonymous):

that was how u do it the work and all?

OpenStudy (anonymous):

that was all the steps worked out. nothing omitted

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

#3 takes two steps subtract 1 from both sides divide by -3 and change the sense of the inequality \[10 < –3x + 1\] \[10-1<-3x\] bet you can finish from there.